how do i solve b squared +3b-4=0 using the box method?

Answer by stanbon(75887)

b^2+3b-4=0
(b+4)(b-1)=0
Either b+4=0 or b-1=0
If b+4=0 then b=-4
If b-1=0 then b=1
Solution: b=-4 or 1.
By the way; what is the "box" method?
Cheers,
Stan H.